OpenStudy (anonymous):

Medals will be given, easy direct variation problem, please help? Multiple choice.

OpenStudy (anonymous):

In a direct variation, y is negative twelve when x is 4. What is y when x equals six?

OpenStudy (anonymous):

Help? c;

OpenStudy (anonymous):

i am not sure y = -12 when x= 4 y +12 = 0, x-4 = 0 y+12 = x-4 put x= 6 then y = -10 its just my assumption

OpenStudy (anonymous):

Okay, thanks, I'll look at Mchilds15's then I'll give medals :)

OpenStudy (joshawad):

y = -10 i think

OpenStudy (anonymous):

:) Hang on, I'll se wat Mchilds15 thinks too

OpenStudy (anonymous):

see*

OpenStudy (mchilds15):

It would be y= -10

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

He answered 1st, so I gave it to him :) Thanks y'all

OpenStudy (anonymous):

its ok we can all give each other medals

OpenStudy (anonymous):

Yay~

OpenStudy (anonymous):

everyone give someone a medal

OpenStudy (joshawad):

no problem

OpenStudy (anonymous):

@isag, can you give gokuson a medal?

OpenStudy (anonymous):

yup......:)

OpenStudy (anonymous):

Okay toodles y'all :)